OpenStudy (anonymous):

how do you solve A= h/2x (b1+b2) for b1

OpenStudy (anonymous):

Is the equation: \[A=\left.\left.\frac{h}{2x}\right(b_1+b_2\right)\] ?

OpenStudy (anonymous):

\[A= \frac{ h }{ 2 }x (b1+b2)\]

OpenStudy (anonymous):

Ah cool. K. So you have: \[\eqalign{ &1.\phantom{space}A=\left(\frac{h}{2}\right)x(b_1+b_2) \\ &2.\phantom{space}\frac{A}{x}=\left(\frac{h}{2}\right)(b_1+b_2) \\ &3.\phantom{space}\frac{2A}{x}=h(b_1+b_2) \\ &4.\phantom{space}\frac{2A}{xh}=b_1+b_2 \\ &5.\phantom{space}\frac{2A}{xh}-b_2=b_1 }\] So therefore, \[b_1=\frac{2A}{xh}-b_2\] Please let me know if there is a step you didn't understand

OpenStudy (anonymous):

Numbered them for conveinience

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

thanks! i get it now!